# SAI3D: Segment Any Instance in 3D Scenes

### 17 Dec 2023 v1

@@ Abstract
Advancements in 3D instance segmentation have traditionally been tethered to the availability of annotated datasets, limiting their application to a narrow spectrum of object categories. Recent efforts have sought to harness vision-language models like CLIP for open-set semantic reasoning, yet these methods struggle to distinguish between objects of the same categories and rely on specific prompts that are not universally applicable. In this paper, we introduce SAI3D, a novel zero-shot 3D instance segmentation approach that synergistically leverages geometric priors and semantic cues derived from Segment Anything Model (SAM). Our method partitions a 3D scene into geometric primitives, which are then progressively merged into 3D instance segmentations that are consistent with the multi-view SAM masks. Moreover, we design a hierarchical region-growing algorithm with a dynamic thresholding mechanism, which largely improves the robustness of fine-grained 3D scene parsing. Empirical evaluations on ScanNet and the more challenging ScanNet++ datasets demonstrate the superiority of our approach. Notably, SAI3D outperforms existing open-vocabulary baselines and even surpasses fully-supervised methods in class-agnostic segmentation on ScanNet++. Our project page is at https://yd-yin.github.io/SAI3D.
@@ 1 Introduction
In this work, we investigate how to better leverage geometric priors and multi-view consistency for fine-grained instance segmentation of intricate 3D scenes. We introduce SAI3D (Segment Any Instance in 3D), which takes advantage of SAM to conduct zero-shot 3D instance segmentation, without training or finetuning with 3D annotations. SAI3D dissects the 3D scene into geometric primitives, and builds a sparse affinity matrix that captures pairwise similarity scores based on the 2D masks generated by SAM. Our approach involves the progressive merging of these primitives using a region-growing algorithm, and an aggregation of votes from all valid images, to obtain a multi-view consistent 3D instance segmentation.
We evaluate our method on ScanNet [4] as well as the recently developed ScanNet++ [50]. Notably, ScanNet++ provides more detailed segmentation masks with rich semantics, thereby offering a more realistic and challenging benchmark for in-the-wild scenarios. For class-agnostic segmentation, our method significantly outperforms SAM3D and other open-vocabulary segmentation baselines on both datasets. Remarkably, when evaluated on ScanNet++, our zero-shot method achieves better results when compared with fully-supervised Mask3D models trained on ScanNet, underlining the efficacy of SAI3D in fine-grained instance segmentation of complex 3D scenes.
@@
  • We present a carefully designed aggregation method of 2D image masks into a coherent 3D segmentations that are consistent across different views.
@@ 2 Related Work
3D semantic segmentation is a long-studied topic, which aims to categorize each point in a given 3D scene with a specific semantic class [32, 33, 24, 47, 22, 9, 14, 46, 37, 31, 51]. 3D instance segmentation extends this by identifying distinct objects within the same semantic category and assigning unique masks to each object instance [3, 29, 43, 36, 45, 14, 40, 42, 5, 11, 13, 26, 44, 16, 21, 41]. ScanNet200 [4] is a standard benchmark used for indoor 3D instance segmentation evaluation, and Mask3D [40] achieves state-of-the-art performance on it using a transformer-based network. Despite its advancements, Mask3D still requires a large amount of 3D annotated data for network training, as previous supervised learning methods. This hampers generalizing the method towards open-world scenarios containing novel objects of unseen categories. Moreover, the annotated 3D data is expensive to collect, and sometimes even impossible due to privacy reasons. In this paper, we focus on zero-shot open-vocabulary 3D segmentation, where no training or finetuning with 3D annotation is needed.
Open-vocabulary 2D image segmentation.
@@
By training on the large-scale web data of image-text pairs, foundation image-language models such as CLIP [34] have achieved impressive performance in aligning image and text in high-dimensional feature space. The following works apply CLIP feature to various zero-shot image tasks, including image captioning [48, 2], object recognition [12], and object detection [27, 10]. More recently, OpenSeg [8] and OV-Seg [25] extend foundation image-language models to semantic image segmentation by learning a semantic-aware pixel-wise embedding. Segment Anything Model (SAM) [19] takes a step further to segment any object, in any image, with user-provided or automatically-generated prompts. SAM has learned a general notion of what objects are, which enables zero-shot generalization to unfamiliar objects and images without requiring additional training. In our work, we benefit from the zero-shot generalization of SAM to produce high-quality 2D masks on multi-view images, and aggregate them into consistent 3D segments using a primitive-based region growing.
Open-vocabulary 3D segmentation.
@@
Inspired by the successful 2D open-vocabulary segmentation models, OpenScene [31] proposes an open-vocabulary 3D semantic segmentation by distilling the CLIP feature onto 3D point clouds. On the other hand, LERF [18] and DFF [20] integrate language within NeRF by optimizing an extra feature field that aligns with the CLIP feature. While these methods enable open-vocabulary querying by text prompts, they cannot distinguish between object instances of the same categories. To solve this issue, OpenMask3D [42] and OpenIns3D [15] leverage the pre-trained Mask3D models for class-agnostic 3D mask proposal generation. While achieving promising instance segmentation results on indoor scenes with similar objects as the training data (ScanNet), we show in our experiments that they fail in complex scenes with fine-grained objects. In this work, we remove the dependency of a pre-trained 3D mask proposal network, and work directly on the geometry-aware 3D primitives.
@@ 3.2 Primitive Affinity
Affinity in a single view. Based on the projected primitive $Q_{i,m}^{\text{2D}}$ and the image segmentation $S_{m}$, we collect the mask labels covered by $Q_{i,m}^{\text{2D}}$ and compute a normalized histogram of the mask labels. This histogram represents a distribution of 2D instance mask labels corresponding to the projected primitive. As shown in Fig. 3, a projected primitive can have multiple 2D instance mask labels, represented in different colors.
We note the mask label distribution of $Q_{i,m}^{\text{2D}}$ as ${\bf h}_{i,m}$, which can be regarded as a directional vector.We compute the affinity score between two primitives projected in the $m$-th image as the cosine similarity between two vectors:
@@
The weight $w_{i,j}^{m}$ is calculated as the product of visibilities of $Q_{i,m}^{\text{2D}}$ and $Q_{j,m}^{\text{2D}}$. Note that $w_{i,j}^{m}=0$ for invalid images where either one of the primitives is not visible.
@@ 3.3 Primitive Merging
Multi-level merging criteria. Given a region represented as a queue of primitives, the top primitive is popped out and used to retrieve the neighboring nodes for growing. In the vanilla region growing algorithm, a node would be added to the region if it shares a high-affinity score with the popped node. However, this pairwise comparison is prone to errors, which would accumulate along with the growing process, as shown in Fig. 4. To solve this problem, we propose a multi-level merging criteria that computes the affinity score in a hierarchical way, where the affinity scores between the candidate node and all the nodes within the region are summed together by weighting them according to the graph distance.
@@
Progressive growing. Another important hyper-parameter of region-growing is the affinity threshold, which is used to determine if two regions should be merged or not. As shown in Fig. 5, the region-growing process is sensitive to the threshold value, and setting a fixed threshold usually results in over-segmentation or under-segmentation. Based on this observation, we propose a progressive growing algorithm, where the growing process is decomposed into multiple stages with threshold values varying from high to low. In this way, we merge small regions with a strict criteria at the beginning, which prevents incorrect merges from accumulating along with the growing process. As the regions gradually grow into large ones that are more reliable, we use a more relaxed criteria to merge them. This dynamic thresholding approach ensures that our merging framework remains sensitive to the evolving certainty of connectivity, thereby enhancing the accuracy of the final segmentation.
@@ 4 Experiment
We evaluate SAI3D on ScanNet++ and ScanNet to demonstrate its effectiveness in 3D instance segmentation. We compare it with leading methods in open-vocabulary segmentation, which, like SAI3D, are designed for zero-shot transfer scenarios. Additionally, we benchmark against the state-of-the-art closed-vocabulary methods that require training on annotated datasets.
@@ 4.1 Experiment Setup
Datasets
ScanNet++ [50] is a very recently released indoor dataset that offers posed RGB-D streams, high-quality 3D geometry captured by advanced laser technology, and comprehensive object annotations. Compared with previous datasets, ScanNet++ features higher-resolution 3D geometry and finer-grained data annotations, especially for long-tail semantics. ScanNet++ poses new challenges that reflect real-world applications, making it ideal for evaluation in terms of open-vocabulary methods. Following Mask3D [40], we also adopt the well-studied dataset ScanNet200 [37], which builds upon the classic ScanNet dataset [4] and extends the semantic annotations to 200 classes.
Evaluation metrics.
We evaluate the numerical results with the widely-used Average Precision score. Following the baselines [42, 40, 38], we report scores at IoU scores of 25 % and 50 % (AP@25, AP@50) and averaged over all overlaps between [50 % and 95 %] at 5 % steps. We adopt two evaluation setups: class-agnostic instance segmentation focuses only on the accuracy of the instance masks themselves, and semantic instance segmentation that also considers their associated semantic labels. We calculated the average score across all semantic categories to obtain the overall performance.
Baselines.
We compare our approach with both open-vocabulary and closed-vocabulary baselines. For open-vocabulary methods, we compare against SAM3D [49] and UnScene3D [38], which both leverage information extracted from large pre-trained models. Notably, SAM3D is very similar to our approach by building upon the automatic mask generation of SAM, but differs in the merging process. Unlike our approach based on 3D primitives, SAM3D projects 2D masks onto 3D and iteratively merges them frame by frame, which overlooks the global geometry properties of the 3D scene. In addition, We also compare with the traditional point grouping methods like HDBSCAN [28] and Felzenszwalb’s algorithm [6], as well as a feature clustering method [30]. For closed-vocabulary comparison, we select Mask3D [40], the state-of-the-art transformer-based method, as our baseline.
@@ 4.2 Results
[…]
Closed-vocabulary
Mask3D [39] ScanNet200 10.2 17.3 24.9
Open-vocabulary
Felzenszwalb et al. [6] - 4.1 9.2 25.3
SAM3D [49] - 7.2 14.2 29.4
Ours - 17.1 31.1 49.5
Table 1: Class-agnostic 3D instance segmentation on ScanNet++ dataset. We compare against both closed-vocabulary and open-vocabulary methods, and report Average Precision scores. Note that Mask3D is trained on ScanNet200 dataset.
Fine-grained 3D segmentation.
Numerical results of the class-agnostic instance segmentation evaluation are shown in Table 1, where our method outperforms prior works in all metrics. Not only did it surpass the unsupervised methods significantly, but it also outperformed the closed-vocabulary baseline trained on ScanNet200. This highlights our method’s ability to handle detailed objects and diverse semantic categories effectively.
Open-vocabulary 3D object querying.
An important application of our method is zero-shot prompt-based instance segmentation. As depicted in Fig. 1 and 7, our approach effectively segments the 3D scene into clean segments, and identifies specific objects with the input prompts. Since our approach generates more detailed and accurate 3D instance segmentation than previous works, we are able to retrieve rare objects like “toilet roll” or “shower gel”.
@@
UnScene3D [38] - 15.9 32.2 58.5
@@
Table 2: Class-agnostic 3D instance segmentation on ScanNet200 dataset. For supervised method Mask3D, we choose two models trained on ScanNet200 and S3DIS, respectively.
Standard 3D segmentation.
Table 2 reports the class-agnostic instance segmentation results on ScanNet200 dataset, where our method significantly outperforms the open-vocabulary baselines. While our approach does not require any training or finetuning on 3D annotation, the supervised method Mask3D suffers from a serious performance drop when transferring from one dataset to another. As can be seen in the table, the performance of Mask3D drops from 60.8 to 31.1 in terms of AP due to the domain gap between ScanNet200 and S3DIS. This indicates that supervised methods tend to overfit the training data, and lack a generalization ability towards open scenes.
@@
Mask3D [39] Supervised Open-vocab 11.7 15.2 17.8 13.4 11.6 9.9
SAM3D [49] Open-vocab Open-vocab 7.4 11.2 16.2 6.7 8.0 7.6
Ours Open-vocab Open-vocab 9.6 14.7 19.0 9.2 10.5 9.1
Table 3: Numerical results of semantic instance segmentation on ScanNet200 dataset. We report the AP scores averaged on all classes except “wall” and “floor”, similar to the baselines.
We also evaluate the semantic instance segmentation on ScanNet200 and report the quantitative results in Table 3. We follow OpenMask3D [42] to generate semantic labels, which extract the pre-trained per-point OpenScene features [31] and compute the cosine similarity between text-mask pairs. And we keep the mask proposal stage of Mask3D supervised, but modify the semantic label assignment to match the open-vocabulary baselines in order to focus on evaluating the quality of mask proposals. As shown in the table, our method clearly results in superior performance than the open-vocabulary baseline, but lags behind the closed-vocabulary Mask3D. Upon closer examination, we notice that our method shows weaker performance on frequently occurring semantics (Head), but the gap narrows with long-tail labels (Tail). This observation highlights the strength of our method in zero-shot generalization, better at handling diverse and less common labels.

### 24 Mar 2024 v2

@@ Abstract
Advancements in 3D instance segmentation have traditionally been tethered to the availability of annotated datasets, limiting their application to a narrow spectrum of object categories. Recent efforts have sought to harness vision-language models like CLIP for open-set semantic reasoning, yet these methods struggle to distinguish between objects of the same categories and rely on specific prompts that are not universally applicable. In this paper, we introduce SAI3D, a novel zero-shot 3D instance segmentation approach that synergistically leverages geometric priors and semantic cues derived from Segment Anything Model (SAM). Our method partitions a 3D scene into geometric primitives, which are then progressively merged into 3D instance segmentations that are consistent with the multi-view SAM masks. Moreover, we design a hierarchical region-growing algorithm with a dynamic thresholding mechanism, which largely improves the robustness of fine-grained 3D scene parsing. Empirical evaluations on ScanNet, Matterport3D and the more challenging ScanNet++ datasets demonstrate the superiority of our approach. Notably, SAI3D outperforms existing open-vocabulary baselines and even surpasses fully-supervised methods in class-agnostic segmentation on ScanNet++. Our project page is at https://yd-yin.github.io/SAI3D.
@@ 1 Introduction
In this work, we investigate how to better leverage geometric priors and multi-view consistency for fine-grained instance segmentation of intricate 3D scenes. We introduce SAI3D (Segment Any Instance in 3D Scenes), which takes advantage of SAM to conduct zero-shot 3D instance segmentation, without training or finetuning with 3D annotations. SAI3D dissects the 3D scene into geometric primitives, and builds a sparse affinity matrix that captures pairwise similarity scores based on the 2D masks generated by SAM. Our approach involves the progressive merging of these primitives using a region-growing algorithm, and an aggregation of votes from all valid images, to obtain a multi-view consistent 3D instance segmentation.
We evaluate our method on ScanNetV2 [5], ScanNet200 [40], Matterport3D [2] as well as the recently developed ScanNet++ [54]. Notably, ScanNet++ provides more detailed segmentation masks with rich semantics, thereby offering a more realistic and challenging benchmark for in-the-wild scenarios. For class-agnostic segmentation, our method significantly outperforms SAM3D and other open-vocabulary segmentation baselines on both datasets. Remarkably, when evaluated on ScanNet++, our zero-shot method achieves better results when compared with fully-supervised Mask3D models trained on ScanNet, underlining the efficacy of SAI3D in fine-grained instance segmentation of complex 3D scenes.
@@
  • We present a carefully designed aggregation method of 2D image masks into coherent 3D segmentations that are consistent across different views.
@@ 2 Related Work
Closed-vocabulary 3D segmentation. 3D semantic segmentation is a long-studied topic, which aims to categorize each point in a given 3D scene with a specific semantic class [35, 36, 25, 50, 23, 10, 15, 49, 40, 34, 55]. 3D instance segmentation extends this by identifying distinct objects within the same semantic category and assigning unique masks to each object instance [4, 31, 46, 39, 48, 15, 43, 45, 6, 12, 14, 27, 47, 17, 22, 44]. ScanNet200 [5] is a standard benchmark used for indoor 3D instance segmentation evaluation, and Mask3D [43] achieves state-of-the-art performance on it using a transformer-based network. Despite its advancements, Mask3D still requires a large amount of 3D annotated data for network training, as previous supervised learning methods. This hampers generalizing the method towards open-world scenarios containing novel objects of unseen categories. Moreover, the annotated 3D data is expensive to collect, and sometimes even impossible due to privacy reasons. In this paper, we focus on zero-shot open-vocabulary 3D segmentation, where no training or finetuning with 3D annotation is needed.
@@
Open-vocabulary 2D image segmentation. By training on the large-scale web data of image-text pairs, foundation image-language models such as CLIP [37] have achieved impressive performance in aligning image and text in high-dimensional feature space. The following works apply CLIP feature to various zero-shot image tasks, including image captioning [51, 3], object recognition [13], and object detection [28, 11]. More recently, OpenSeg [9] and OV-Seg [26] extend foundation image-language models to semantic image segmentation by learning a semantic-aware pixel-wise embedding. Segment Anything Model (SAM) [20] takes a step further to segment any object, in any image, with user-provided or automatically-generated prompts. SAM has learned a general notion of what objects are, which enables zero-shot generalization to unfamiliar objects and images without requiring additional training. In our work, we benefit from the zero-shot generalization of SAM to produce high-quality 2D masks on multi-view images, and aggregate them into consistent 3D segments using a primitive-based region growing.
@@
Open-vocabulary 3D segmentation. Inspired by the successful 2D open-vocabulary segmentation models, OpenScene [34] proposes an open-vocabulary 3D semantic segmentation by distilling the CLIP feature onto 3D point clouds. On the other hand, LERF [19] and DFF [21] integrate language within NeRF by optimizing an extra feature field that aligns with the CLIP feature. While these methods enable open-vocabulary querying by text prompts, they cannot distinguish between object instances of the same categories. To solve this issue, OpenMask3D [45] and OpenIns3D [16] leverage the pre-trained Mask3D models for class-agnostic 3D mask proposal generation. While achieving promising instance segmentation results on indoor scenes with similar objects as the training data (ScanNet), we show in our experiments that they fail in complex scenes with fine-grained objects. Concurrent work, MaskClustering [52] and Open3DIS [32], both utilize 2D foundation models to obtain segmentation masks, subsequently aggregating these masks to construct 3D representations. In this work, we integrate both 3D geometric priors and semantic clues from multi-view 2D masks from SAM model for fine-grained 3D instance segmentation.
@@ 3.2 Primitive Affinity
Affinity in a single view. Based on the projected primitive $Q_{i,m}^{\text{2D}}$ and the image segmentation $S_{m}$, we collect the mask labels covered by $Q_{i,m}^{\text{2D}}$ and compute a normalized histogram of the mask labels as a vector, denoted as $\mathbf{h}_{i,m}$. This histogram represents a distribution of 2D instance mask labels corresponding to the projected primitive. As shown in Fig. 3, a projected primitive can have multiple 2D instance mask labels, represented in different colors.
We compute the affinity score between two primitives projected in the $m$-th image as the cosine similarity between two vectors:
@@
The weight $w_{i,j}^{m}$ is calculated as the product of visibilities of $Q_{i,m}^{\text{2D}}$ and $Q_{j,m}^{\text{2D}}$ as Eq. 4
$\footnotesize w_{i,j}^{m}=\frac{\sum_{p\in Q_{i}}\mathbbm{1}\left(\text{Valid}(p,S_{m})\right)}{|Q_{i}|}\frac{\sum_{p\in Q_{j}}\mathbbm{1}\left(\text{Valid}(p,S_{m})\right)}{|Q_{j}|}$ (4)
where “Valid” function indicates if the projected point $p$ is visible in the 2D segmentation mask $S_{m}$. Note that $w_{i,j}^{m}=0$ for invalid images where either one of the primitives is not visible.
@@ 3.3 Primitive Merging
Multi-level merging criteria. Given a region represented as a queue of primitives, the top primitive is popped out and used to retrieve the neighboring nodes for growing. In the vanilla region growing algorithm, a node would be added to the region if it shares a high-affinity score with the popped node. However, this pairwise comparison is prone to errors, which would accumulate along with the growing process, as shown in Fig. 4. To solve this problem, we propose a multi-level merging criteria that compute the affinity score in a hierarchical way, where the affinity scores between the candidate node and all the nodes within the region are summed together by weighting them according to the graph distance.
@@
Progressive growing. Another important hyper-parameter of region-growing is the affinity threshold, which is used to determine if two regions should be merged or not. As shown in Fig. 5, the region-growing process is sensitive to the threshold value, and setting a fixed threshold usually results in over-segmentation or under-segmentation. Based on this observation, we propose a progressive growing algorithm, where the growing process is decomposed into multiple stages with threshold values varying from high to low. In this way, we merge small regions with a strict criterion at the beginning, which prevents incorrect merges from accumulating along with the growing process. As the regions gradually grow into large ones that are more reliable, we use a more relaxed criterion to merge them. This dynamic thresholding approach ensures that our merging framework remains sensitive to the evolving certainty of connectivity, thereby enhancing the accuracy of the final segmentation.
@@ 4 Experiment
We evaluate SAI3D on multiple datasets to demonstrate its effectiveness in 3D instance segmentation. We compare it with leading methods in open-vocabulary segmentation, which, like SAI3D, are designed for zero-shot transfer scenarios. Additionally, we benchmark against the state-of-the-art closed-vocabulary methods that require training on annotated datasets.
@@ 4.1 Experiment Setup
Datasets. ScanNet++ [54] is a very recently released indoor dataset that offers posed RGB-D streams, high-quality 3D geometry captured by advanced laser technology, and comprehensive object annotations. Compared with previous datasets, ScanNet++ features higher-resolution 3D geometry and finer-grained data annotations, especially for long-tail semantics, which poses new challenges that reflect real-world applications. To better validate the robustness of our method, we also incorporate the well-studied ScanNetV2 [5], ScanNet200 [40] (with 200 semantic classes) and Matterport3D[2] datasets. Please see supplementary for more details.
Evaluation metrics. We evaluate the numerical results with the widely-used Average Precision score. Following the baselines [45, 43, 41], we report scores at IoU scores of 25 % and 50 % (AP@25, AP@50) and averaged over all overlaps between [50 % and 95 %] at 5 % steps. We adopt two evaluation setups: class-agnostic instance segmentation focuses only on the accuracy of the instance masks themselves, and semantic instance segmentation that also considers their associated semantic labels. We calculated the average score across all semantic categories to obtain the overall performance.
Baselines. We compare our approach with both closed-vocabulary and open-vocabulary baselines. Mask3D is the state-of-the-art, transformer-based method, supervised with training data. For open-vocabulary methods, we compare with the recent SAM3D [53], UnScene3D [41], and OVIR-3D [29]. Notably, SAM3D is very similar to our approach by building upon the automatic mask generation of SAM, but differs in the merging process. Unlike our approach based on 3D primitives, SAM3D projects 2D masks onto 3D and iteratively merges them frame by frame, which overlooks the global geometry properties of the 3D scene. OpenMask3D [45] is built upon Mask3D with supervised mask proposals and open-vocabulary semantic assignment. In addition, We also compare with the traditional point grouping methods like HDBSCAN [30] and Felzenszwalb’s algorithm [7], as well as a feature clustering method [33].
@@ 4.2 Results
Table 1: Class-agnostic 3D instance segmentation on ScanNet++ dataset. We compare against both closed-vocabulary and open-vocabulary methods, and report average precision scores. Note that Mask3D is trained on ScanNetV2 dataset.
[…]
Closed-vocabulary
Mask3D [42] ScanNetV2 9.9 17.3 25.8
Open-vocabulary
Felzenszwalb et al. [7] - 4.1 9.2 25.3
SAM3D [53] - 7.2 14.2 29.4
Ours - 17.1 31.1 49.5
Fine-grained 3D segmentation. Table 1 reports the numerical results for class-agnostic instance segmentation on ScanNet++ dataset. Our method outperforms prior works in all evaluation metrics. Not only did it surpass the unsupervised methods significantly, but it also outperformed the closed-vocabulary baseline trained on ScanNetV2. This highlights our method’s ability to handle detailed objects and diverse semantic categories effectively.
Open-vocabulary 3D object querying. An important application of our method is zero-shot prompt-based instance segmentation. As depicted in Fig. 1 and 7, our approach effectively segments the 3D scene into clean segments, and identifies specific objects with the input prompts. Since our approach generates more detailed and accurate 3D instance segmentation than previous works, we are able to retrieve rare objects like “toilet roll” or “shower gel”.
@@
Table 2: Class-agnostic 3D instance segmentation on ScanNetV2 dataset.
@@
UnScene3D [41] ***Due to the unavailability of the code, we follow its experiment settings to evaluate on ScanNetV2 dataset. - 15.9 32.2 58.5
@@
Standard 3D segmentation. Table 2 reports the class-agnostic instance segmentation results on ScanNetV2 dataset. Our method significantly outperforms the open-vocabulary baselines that do not require any training or finetuning on 3D annotation. As seen from the table, the supervised method Mask3D suffers from a performance drop when trained on a different dataset, which indicates that supervised methods tend to overfit the training data, and lack a generalization ability toward open scenes.
Table 3: Semantic instance segmentation on ScanNet200 dataset.
@@
Sup. mask + Open-vocab. semantic
OpenMask3D [45] 15.4 19.9 23.1 17.1 14.1 14.9
Open-vocab. mask + Open-vocab. semantic
OVIR-3D [29]The numbers are different from those in the original paper, since the original paper adopts a different evaluation metric namely mAP for information retrieval. 9.3 18.7 25.0 9.8 9.4 8.5
SAM3D [53] 9.8 15.2 20.7 9.2 8.3 12.3
Ours 12.7 18.8 24.1 12.1 10.4 16.2
Following OpenMask3D [45], semantic instance segmentation is evaluated on ScanNet200 dataset, with numerical results in Table 3. We adopt the pipeline of OpenMask3D [45] to assign semantic labels for our method. As shown in the table, our method clearly results in superior performance than the open-vocabulary baselines, but lags behind OpenMask3D that relies on the supervised mask proposals. Upon closer examination, we notice that our method shows weaker performance on frequently occurring semantics (Head), but outperforms on long-tail labels (Tail). This observation highlights the strength of our method in zero-shot generalization, better at handling diverse and less common labels.
To further evaluate the generalizability, we conduct experiments on Matterport3D dataset for both class-agnostic and semantic instance segmentation tasks. As shown in Table 4, our zero-shot method consistently outperforms other baselines, while OpenMask3D suffers a significant performance drop since it is trained on ScanNet200 dataset.
Table 4: Class-agnostic and semantic instance segmentation on Matterport3D dataset.
Method Class-agnostic Semantic
[…]
OpenMask3D [45] 15.3 28.3 43.3 7.7 13.9 20.3
OVIR-3D [29] 6.6 15.6 28.3 5.8 13.8 22.4
SAM3D [53] 10.1 19.4 36.1 4.4 7.3 11.3
Ours 21.5 38.3 59.1 8.9 15.3 20.9